OpenStudy (anonymous):

5. A royal flush consists of an ace, king, queen, jack, and 10 all in the same suit. If 7 cards are dealt at random from a standard deck, determine the probability of getting a) a royal flush in spades. b) a royal flush in any suit.

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

So first how many cards are in a deck?

OpenStudy (anonymous):

@iowastate

OpenStudy (anonymous):

52

OpenStudy (anonymous):

okay now how many different suits are there?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

there are a couple of ways to do this you have 5 cards to choose and 52 total probability the first card is one of those 5 is \(\frac{5}{52}\) now there are 4 cards to choose from and 51 cards, next probability is \(\frac{4}{51}\) and so on multiply to get your answer

OpenStudy (ybarrap):

 There are \({52 \choose 7}\) possible ways to deal the card. You want 1) Ways of grouping 5 cards that make the royal flush in a single group of 5 times ways of grouping the remaining 47 cards in a group of two:\({5\choose 5}\times {47\choose2}\) 2) Number of suits that will form a royal flush:\(4\choose 1\) 3) Possible ways to deal cards:\({52 \choose 7}\) But you are looking only for a royal flush that is a \(spade\), which is just 1/4 of the previous result because there are 4 suits and thus four ways to win. $$ \large \cfrac{1}{4}\cfrac{{4\choose 1}{5\choose 5}{47\choose2}}{{52\choose 7}} $$